# Discrete Audio Tokens: More Than a Survey!

### v2

@@ 1 Introduction
Recent studies have introduced a variety of tokenization methods, often grouped into two main categories: acoustic and semantic [^1] It is important to clarify that the term “semantic” in the speech context does not align with its conventional linguistic meaning. In the speech context, these discrete tokens are more accurately described as phonetic units (Sicherman & Adi, 2023; Choi et al., 2024) and typically do not carry semantic content (Arora et al., 2025). In this paper, to maintain consistency across different domains (speech, audio, music) and with established terminology such as “semantic distillation,” we consistently use the term “semantic.”  (Borsos et al., 2023b; Zhang et al., 2024a; Har-Tuv et al., 2025; Guo et al., 2025b). Acoustic tokens are typically learned through encoder-decoder architectures optimized for waveform reconstruction (Zeghidour et al., 2021; Défossez et al., 2023; Kumar et al., 2023; Yang et al., 2023a). Semantic tokens are derived from pretrained self-supervised learning (SSL) models (Lakhotia et al., 2021; Mousavi et al., 2024b) or encoders trained in a supervised manner (Du et al., 2024b) designed to capture phonetic or linguistic content for discriminative tasks such as speech recognition and translation. Some recent approaches aim to combine both types, introducing hybrid tokenizers (Zhang et al., 2024a; Défossez et al., 2024) that balance acoustic and phonetic properties.
@@ 2.1 Overall architecture
  • An encoder that converts the input waveform ${}{x}$ into a sequence of frame-wise embeddings ${Z}=\{{z}_{t}\}_{t=1}^{T}$ using an encoder function $f_{e}$, such that, ${Z}=f_{e}({x})$, where each ${z}_{t}\in\mathbb{R}^{D}$ is a continuous embedding at time step $t$.
@@
  • A decoder that reconstructs the waveform from the sequence of quantized embeddings using a decoder function $f_{d}$, i.e., $\hat{{x}}=f_{d}(\hat{{z}})$, where $\hat{{x}}$ is the reconstructed waveform.
@@ 2.2.1 Quantization Algorithm
SVQ uses a single codebook for quantization, where each frame-wise embedding is mapped to a single code, unlike RVQ, which uses multiple codes. SVQ can be viewed as a simplified form of RVQ with a single codebook (i.e., $M{=}1$), without iterative residual refinement, similar to VQ-VAE (Gârbacea et al., 2019). It has gained popularity due to the architectural complexity introduced by multiple codebooks in acoustic language models, such as managing multiple codebook streams. SVQ, by contrast, is simpler and particularly useful for training acoustic language models. To compensate for the potential loss of information caused by using a single codebook, some SVQ-based codec models adopt larger codebook sizes. Examples of SVQ-based codecs include BigCodec (Xin et al., 2024), TS3-Codec (Wu et al., 2024d), WavTokenizer (Ji et al., 2024c).
@@
[…]
@@
Unlike traditional vector quantization, FSQ maps each dimension of a feature vector to a fixed set of scalar values (Mentzer et al., 2024). Specifically, each feature value $z_{t}$ is first squashed into the range $[-1,1]$ using a non-linear function such as $\tanh$, and then quantized into a scalar latent space by computing $\text{round}(z_{t}\cdot S)/S$, where $S$ is a hyperparameter controlling the quantization resolution. This procedure results in $12S+$ 1 distinct scalar values per dimension, ensuring uniform coverage of the latent space. FSQ has been adopted in various recent models. SQ-Codec (Yang et al., 2024d) achieves this by creating a scalar latent space, while Spectral Codecs (Langman et al., 2024) use FSQ to encode mel-spectrogram features into a flat codebook. FSQ is often used with diffusion models for high-quality audio generation. HARP-Net (Petermann et al., 2021) similarly applies FSQ but directly maps bottleneck features i.e., single scalar values rather than vectors, to a set of learned scalar bins. Unlike other approaches, HARP-Net maintains the original input frame rate (44.1 kHz) by avoiding temporal decimation, instead expanding the feature dimension in intermediate layers before collapsing to scalar quantization. FocalCodec (Della Libera et al., 2025) instead uses a variant of FSQ called Binary Spherical Quantization (BSQ), which relies on two scalar values.
@@ 2.4.1 Training Strategies
Training strategies for audio tokenizers can be categorized into two broad approaches: separate (post-training) and joint (end-to-end training). These differ in how the encoder, quantizer, and decoder modules are optimized in relation to each other. In both cases, the encoder may be randomly initialized or initialized using a pretrained model, such a SSL model (e.g., wav2vec 2.0 (Baevski et al., 2020), HuBERT (Hsu et al., 2021), WavLM (Chen et al., 2022)) or an ASR model (Radford et al., 2023)).
@@ 2.4.2 Main Training Objectives
where $\text{sg}[\cdot]$ denotes the stop-gradient operator. This loss penalizes discrepancies between the encoder outputs and their corresponding quantized embeddings while ensuring that gradients do not update the codebook entries directly. Modern approaches often replace this loss with Exponential Moving Average (EMA) updates for the codebook, which improve training stability and mitigate codebook collapse.
@@
where $D_{k}(\cdot)$ denotes the output of the $k$-th discriminator. To improve the quality of the generated audio, multiple discriminators are commonly used, including multi-resolution and multi-scale short-time Fourier transform (STFT) discriminators from neural vocoder research (Kong et al., 2020; Défossez et al., 2023; Zeghidour et al., 2021). Perceptual losses are applied using a variety of discriminators, including multi-scale STFT-based (MS-STFT), multi-period (MPD), and multi-scale (MSD) variants.
@@
Masked prediction loss is commonly used in tokenizers where the encoder and decoder are trained separately. The encoder is trained to predict masked portions of the input sequence, typically capturing phonetic information rather than reconstructing the full waveform. Following the masked language modeling (MLM) paradigm (Devlin et al., 2019), a portion of the input is randomly masked, and the model is optimized to predict the masked frames from the surrounding context. This approach is widely used in speech pretraining models such as HuBERT (Hsu et al., 2021) and WavLM (Chen et al., 2022) and is adopted in several semantic tokenizer designs (Lakhotia et al., 2021; Mousavi et al., 2024b).
@@ 2.4.3 Auxiliary Components
Disentanglement methods separate different speech attributes into distinct representations, reducing redundancy while allowing independent control over acoustic properties and simplifying downstream tasks. One line of disentanglement focuses on separating speech and background audio in the codec embedding space, enabling better bitrate, entropy control (Yang et al., 2021) or speech enhancement (Omran et al., 2023). Another line separates the conceptual and fundamental components of speech. Early attempts obtained efficient and low-bitrate speech coding through speaker and phoneme disentanglement, utilizing separate training (Polyak et al., 2021) or joint training (Jiang et al., 2023). More recently, TiCodec (Ren et al., 2024b) minimizes token usage by separately quantizing time-invariant global embeddings (e.g., timbre) and time-varying features (e.g., phonetic information). FACodec (Ju et al., 2024) decomposes speech into subspaces such as content, prosody, timbre, and acoustic details through supervised techniques. The timbre extractor in FACodec is optimized with a speaker classification loss, while distinct RVQ modules process other components before supervised decoupling. LSCodec (Guo et al., 2025a) introduces a low-bitrate, speaker-decoupled speech codec using a three-stage training framework with speaker perturbation. A VQ layer is applied after a VAE that disentangles speaker attributes in a continuous space, followed by training a token vocoder on the quantized codes. Unlike most acoustic tokens that redundantly encode speaker timbre across time steps, LSCodec minimizes this inefficiency by isolating timbre from content and prosody. SoCodec (Guo et al., 2024a) employs multi-stream phonetic sequences and ordered product quantization to encode speech into phonetic and time-variant token sequences using HuBERT as a pretrained SSL model. An ECAPA-TDNN-based encoder (Desplanques et al., 2020) extracts an utterance-level global embedding to retain time-invariant information, such as speaker identity, global speaking style, and acoustic environment. SD-Codec (Bie et al., 2025) integrates audio coding with source separation by assigning different audio domains (such as speech, music, and sound effects) to distinct codebooks using multiple parallel RVQ modules.
@@ 3 Benchmark Evaluation
A summary of all tokenizers included in the benchmark evaluation is provided in Table 2. We utilize pre-trained checkpoints released by the original authors. An overview of our benchmark evaluation pipeline is illustrated in Figure 1.
@@ 3.1 Evaluation for Reconstructed Audio Quality and Complexity
Similar to DAC, WavTokenizer is not explicitly optimized with a time-domain waveform reconstruction loss, resulting in poor SDR and SI-SNR scores. However, its perceptual metrics (VISQOL and SingMOS) remain relatively strong. This further highlights the limitations of time-domain metrics in evaluating token-based representations, as previously observed in the speech reconstruction results. SQ-Codec, the only fully out-of-domain model (trained only on speech data), gives the poorest performance in SDR and SI-SNR metrics. Nevertheless, its perceptual metrics remain comparable to in-domain models at certain bitrates, indicating that despite limited signal fidelity, key perceptual characteristics can still be preserved.
@@ 3.2 Downstream Evaluation
Audio and Music Tasks. For general audio and music tasks, EnCodec consistently outperforms other tokenizers across all bitrates and domains, while DAC lags behind. Although DAC is known for strong perceptual quality, its signal-level fidelity is generally lower, which likely impacts its separation performance. The difficulty of these tasks is evident from the SI-SDR of the unprocessed mixtures, for example, approximately -16 dB for general audio and -7.7 dB for music. Even the best-performing model (EnCodec at medium bitrate) only reaches about -7 dB SI-SDR for audio and -5.7 dB for music. While high-bitrate settings have proven to be challenging for downstream tasks, they perform particularly poorly in music separation, emphasizing that increasing bitrate alone does not improve separation quality and may even degrade performance. This may be due to the inherently polyphonic and less sparse nature of music (in contrast to speech and general audio), which results in highly overlapping sources that are harder to disentangle from detailed but semantically entangled representations.
@@
Semantic tokenizers are generally more robust, especially in low-resource settings, but still fall short of continuous representations when data is limited. Training downstream models with semantic or semantically-distilled tokenizers tends to be more stable and reliable compared to acoustic tokenizers, which often require larger datasets and more careful model scaling. Overall, discrete tokenizers are more sensitive to architectural choices and hyperparameters of the downstream head, whereas continuous features typically yield more consistent performance across configurations. Therefore, careful tuning and appropriate scaling of both data and model architecture are crucial for effectively leveraging discrete representations. While discrete tokens offer advantages in efficiency and modularity, continuous representations still lead in overall performance. Bridging this gap is essential for the successful integration of audio tokens into future multimodal language models.

### v3

@@ 1 Introduction
Recent studies have introduced a variety of tokenization methods, often grouped into two main categories: acoustic and semantic [^1] It is important to clarify that the term “semantic” in the speech context does not align with its conventional linguistic meaning. In the speech context, these discrete tokens are more accurately described as phonetic units (Sicherman & Adi, 2023; Choi et al., 2024) and typically do not carry semantic content (Arora et al., 2025). In this paper, to maintain consistency across different domains (speech, audio, music) and with established terminology such as “semantic distillation,” we consistently use the term “semantic.”  (Borsos et al., 2023b; Zhang et al., 2024a; Har-Tuv et al., 2025; Guo et al., 2025b). Acoustic tokens are typically learned through encoder-decoder architectures optimized for waveform reconstruction (Zeghidour et al., 2021; Défossez et al., 2023; Kumar et al., 2023; Yang et al., 2023a). Semantic tokens are derived from pretrained self-supervised learning (SSL) models (Lakhotia et al., 2021; Mousavi et al., 2024b), which are trained on raw audio without labels by solving proxy tasks (e.g., masked prediction) to learn transferable representations to downstream tasks, or encoders trained in a supervised manner (Du et al., 2024b) designed to capture phonetic or linguistic content for discriminative tasks such as speech recognition and translation. Some recent approaches aim to combine both types, introducing hybrid tokenizers (Zhang et al., 2024a; Défossez et al., 2024) that balance acoustic and phonetic properties.
@@ 2.1 Overall architecture
  • An encoder that converts the input waveform x{}{x} into a sequence of frame-wise embeddings ${Z}=\{{z}_{t}\}_{t=1}^{T}$ using an encoder function $f_{e}$, such that, ${Z}=f_{e}({x})$, where each ${z}_{t}\in\mathbb{R}^{D}$ is a continuous embedding at time step $t$.
@@
[…]
@@ 2.2.1 Quantization Algorithm
SVQ uses a single codebook for quantization, where each frame-wise embedding is mapped to a single code, unlike RVQ, which uses multiple codes. SVQ can be viewed as a simplified form of RVQ with a single codebook (i.e., M=1M{=}1), without iterative residual refinement, similar to VQ-VAE (Gârbacea et al., 2019). It has gained popularity due to the architectural complexity introduced by multiple codebooks in acoustic language models, such as managing multiple codebook streams. SVQ, by contrast, is simpler and particularly useful for training acoustic language models. To compensate for the potential loss of information caused by using a single codebook, some SVQ-based codec models adopt larger codebook sizes. Examples of SVQ-based codecs include BigCodec (Xin et al., 2024), TS3-Codec (Wu et al., 2024d), WavTokenizer (Ji et al., 2024c).
@@
where each $z_{t}^{(g)}\in\mathbb{R}^{D/G}$ represents a segment of the input feature, and $\|$ denotes concatenation. Each group is quantized independently using a separate RVQ module, producing a group-wise quantized embedding $\hat{z}_{t}^{(g)}$. The final quantized vector is formed by concatenating the quantized outputs from all groups:
@@
Unlike traditional vector quantization, FSQ maps each dimension of a feature vector to a fixed set of scalar values (Mentzer et al., 2024). Specifically, each feature value $z_{t}$ is first squashed into the range [1,1][-1,1] using a non-linear function such as $\tanh$, and then quantized into a scalar latent space by computing $\text{round}(z_{t}\cdot S)/S$, where $S$ is a hyperparameter controlling the quantization resolution. This procedure results in $2S+1$ distinct scalar values per dimension, ensuring uniform coverage of the latent space. FSQ has been adopted in various recent models. SQ-Codec (Yang et al., 2024d) achieves this by creating a scalar latent space, while Spectral Codecs (Langman et al., 2024) use FSQ to encode mel-spectrogram features into a flat codebook. FSQ is often used with diffusion models for high-quality audio generation. HARP-Net (Petermann et al., 2021) similarly applies FSQ but directly maps bottleneck features i.e., single scalar values rather than vectors, to a set of learned scalar bins. Unlike other approaches, HARP-Net maintains the original input frame rate (44.1 kHz) by avoiding temporal decimation, instead expanding the feature dimension in intermediate layers before collapsing to scalar quantization. FocalCodec (Della Libera et al., 2025) instead uses a variant of FSQ called Binary Spherical Quantization (BSQ), which relies on two scalar values.
@@ 2.4.1 Training Strategies
Training strategies for audio tokenizers can be categorized into two broad approaches: separate (post-training) and joint (end-to-end training). These differ in how the encoder, quantizer, and decoder modules are optimized in relation to each other. In both cases, the encoder may be randomly initialized or initialized using a pretrained model, such an SSL model (e.g., wav2vec 2.0 (Baevski et al., 2020), HuBERT (Hsu et al., 2021), WavLM (Chen et al., 2022)) or an ASR model (Radford et al., 2023)).
@@ 2.4.2 Main Training Objectives
where $\text{sg}[\cdot]$ denotes the stop-gradient operator. This loss penalizes discrepancies between the encoder outputs and their corresponding quantized embeddings while ensuring that gradients do not update the codebook entries directly. Modern approaches often replace this loss with Exponential Moving Average (EMA) updates for the codebook, which improve training stability and mitigate codebook collapse. This collapse occurs when the model selects only a few code vectors, leaving the rest inactive and unupdated (Dhariwal et al., 2020; Kumar et al., 2023). As a result, the effective codebook size decreases, lowering the target bitrate and degrading reconstruction quality. To address this issue, researchers have proposed several strategies to improve codebook utilization. Some models (Dhariwal et al., 2020; Zeghidour et al., 2021) apply codebook expiration, periodically reinitializing inactive code vectors. Others (Kumar et al., 2023; Yang et al., 2024d; Défossez et al., 2023) use factorized quantization and L2 normalization to encourage more balanced usage across the codebook. Techniques like ESC (Gu & Diao, 2024) and NDVQ (Niu et al., 2024)introduce Euclidean normalization or represent codebooks as distributions, using margin-based or probabilistic losses to prevent collapse and promote diverse activation. Additional solutions introduce auxiliary constraints, such as entropy penalties or code balancing losses. For example, Enhanced Residual Vector Quantization (ERVQ) Zheng et al. (2025) uses intra-codebook optimization via online clustering and a code balancing loss to reactivate unused vectors, while inter-codebook optimization minimizes redundancy between adjacent quantizers through a structural similarity loss, enhancing expressiveness and overall utilization.
@@
where $D_{k}(\cdot)$ denotes the output of the $k$-th discriminator. Following VQGAN Esser et al. (2021), the subsequent audio neural vocoders commonly include multiple discriminators with a specific focus on the frequency-level rebuilding to enhance the perceptual quality (Kong et al., 2020; Défossez et al., 2023; Zeghidour et al., 2021). Specifically, these multi-scale discriminators take the stack of multi-resolution or multi-scale complex-valued short-time Fourier transform (STFT) with the real and imaginary parts concatenated as input, e.g., 5 different scales with STFT window lengths of [2048, 1024, 512, 256, 128], for capturing different structures in audio signals.
@@
Masked prediction loss is commonly used in tokenizers where the encoder and decoder are trained separately. The encoder is trained to predict masked portions of the input sequence, typically capturing phonetic information rather than reconstructing the full waveform. Following the masked language modeling (MLM) paradigm (Devlin et al., 2019), a portion of the input is randomly masked, and the model is optimized to predict the masked frames from the surrounding context. Formally, given an input sequence of frame-level features $\mathbf{X}=\{x_{1},x_{2},\dots,x_{T}\}$, a binary mask $\mathbf{M}\in\{0,1\}^{T}$ is applied, where $M_{t}=1$ indicates a masked position. The masked input is denoted as $\mathbf{X}^{\text{mask}}$, where $x_{t}^{\text{mask}}=[\text{MASK}]$ if $M_{t}=1$, and $x_{t}^{\text{mask}}=x_{t}$ otherwise. The encoder processes this masked input to produce latent representations $\mathbf{Z}=f_{e}(\mathbf{X}^{\text{mask}})$, and the model is trained to minimize prediction loss.
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{L}_{\text{MP}}=\sum_{t=1}^{T}M_{t}\cdot\ell(Z_{t},x_{t})$}
where $f_{e}$ is the encoder network and $\ell(\cdot,\cdot)$ is the cross-entropy loss. This approach is widely used in speech pretraining models such as HuBERT (Hsu et al., 2021) and WavLM (Chen et al., 2022) and is adopted in several semantic tokenizer designs (Lakhotia et al., 2021; Mousavi et al., 2024b)
@@ 2.4.3 Auxiliary Components
Disentanglement methods separate different speech attributes into distinct representations, reducing redundancy while allowing independent control over acoustic properties and simplifying downstream tasks. One type of disentanglement in the codec focuses on separating speech and background audio embedding space, enabling better bitrate, entropy control (Yang et al., 2021) or speech enhancement (Omran et al., 2023). Those models normally aim to find a latent space where two ideally orthogonal components $\mathcal{Z}_{1}$ and $\mathcal{Z}_{2}$ can be conveniently separated, $\mathcal{F}(x)=\mathcal{Z}_{1}\otimes\mathcal{Z}_{2}$, with $\otimes$ representing straightforward operations such as splitting along the channel  (Yang et al., 2021).
Another type separates the conceptual and fundamental components of speech, where each component is typically extracted by its specific encoder. $\mathcal{Z}_{k}=\mathcal{F}_{k}(x)$. Early attempts obtained efficient and low-bitrate speech coding through speaker and phoneme disentanglement, utilizing separate training (Polyak et al., 2021) or joint training (Jiang et al., 2023). More recently, TiCodec (Ren et al., 2024b) minimizes token usage by separately quantizing time-invariant global embeddings (e.g., timbre) and time-varying features (e.g., phonetic information). FACodec (Ju et al., 2024) decomposes speech into subspaces such as content, prosody, timbre, and acoustic details through supervised techniques. The timbre extractor in FACodec is optimized with a speaker classification loss, while distinct RVQ modules process other components before supervised decoupling. LSCodec (Guo et al., 2025a) introduces a low-bitrate, speaker-decoupled speech codec using a three-stage training framework with speaker perturbation. A VQ layer is applied after a VAE that disentangles speaker attributes in a continuous space, followed by training a token vocoder on the quantized codes. Unlike most acoustic tokens that redundantly encode speaker timbre across time steps, LSCodec minimizes this inefficiency by isolating timbre from content and prosody. SoCodec (Guo et al., 2024a) employs multi-stream phonetic sequences and ordered product quantization to encode speech into phonetic and time-variant token sequences using HuBERT as a pretrained SSL model. An ECAPA-TDNN-based encoder (Desplanques et al., 2020) extracts an utterance-level global embedding to retain time-invariant information, such as speaker identity, global speaking style, and acoustic environment. SD-Codec (Bie et al., 2025) integrates audio coding with source separation by assigning different audio domains (such as speech, music, and sound effects) to distinct codebooks using multiple parallel RVQ modules.
@@ 3 Benchmark Evaluation
A summary of all tokenizers included in the benchmark evaluation is provided in Table 2. We select these tokenizers based on several factors: (1) We prioritize open-source models with accessible checkpoints and code to ensure reproducibility; (2) we include tokenizers representing a diverse range of quantization strategies—including RVQ (EnCodec, DAC), SVQ (WavTokenizer), FSQ (SQ-Codec), KMeans (Discrete WavLM), and semantically distilled methods (SpeechTokenizer, Mimi); and (3) we aim to cover multiple domains such as speech, music, and general audio, favoring multi-domain tokenizers where available. The impact of single- versus multi-domain tokenizers is further discussed in Section 4. We utilize pre-trained checkpoints released by the original authors. An overview of our benchmark evaluation pipeline is illustrated in Figure 1.
@@ 3.1 Evaluation for Reconstructed Audio Quality and Complexity
Similar to DAC, WavTokenizer is not explicitly optimized with a time-domain waveform reconstruction loss, resulting in poor SDR and SI-SNR scores. However, its perceptual metrics (VISQOL and SingMOS) remain relatively strong. This further highlights the limitations of time-domain metrics in evaluating token-based representations, as previously observed in the speech reconstruction results. Among the two WavTokenizer variants, WT-S-24 and WT-SMA-24, only WT-S-24 is fully out-of-domain, having been trained exclusively on speech data. Both models exhibit poor performance across objective metrics such as SDR and SI-SNR, as well as perceptual metrics (VISQOL and SingMOS). This degradation stems from a combination of factors: the domain mismatch in WT-S-24, the use of the lowest bitrate among all models, and the absence of explicit waveform reconstruction losses during training.
@@ 3.2 Downstream Evaluation
Audio and Music Tasks. For general audio and music tasks, EnCodec consistently outperforms other tokenizers across all bitrates and domains, while DAC lags behind. Although DAC is known for strong perceptual quality, its signal-level fidelity is generally lower, which likely impacts its separation performance. The difficulty of these tasks is evident from the SI-SDR of the unprocessed mixtures, for example, approximately -16 dB for general audio and -7.7 dB for music. Even the best-performing model (EnCodec at medium bitrate) only reaches about -7 dB SI-SDR for audio and -5.7 dB for music. We report the performance in terms of SI-SDR improvement (“SI-SDRi”), not absolute SI-SDR. Thus, the reason we also provide the performance on “unprocessed” mixtures. For instance, for general audio, we report an improvement of 9.53 dB over the mixture (-16.5 dB). In absolute value, this means that the resulting predictions yield an average of -7-7 dB performance. While high-bitrate settings have proven to be challenging for downstream tasks, they perform particularly poorly in music separation, emphasizing that increasing bitrate alone does not improve separation quality and may even degrade performance. This may be due to the inherently polyphonic and less sparse nature of music (in contrast to speech and general audio), which results in highly overlapping sources that are harder to disentangle from detailed but semantically entangled representations.
@@
Semantic tokenizers (e.g., Discrete WavLM) are generally more robust, especially in low-resource settings, but still fall short of continuous representations when data is limited. Training downstream models with semantic (Discrete WavLM) or semantically distilled tokenizers (Mimi and SpeechTokenizer) tends to be more stable and reliable compared to acoustic tokenizers (EnCodec, DAC, WavTokenizer, and SQ-Codec), which often require larger datasets and more careful model scaling. Overall, discrete tokenizers are more sensitive to architectural choices and hyperparameters of the downstream head, whereas continuous features typically yield more consistent performance across configurations. Therefore, careful tuning and appropriate scaling of both data and model architecture are crucial for effectively leveraging discrete representations. While discrete tokens offer advantages in efficiency and modularity, continuous representations still lead in overall performance. Bridging this gap is essential for the successful integration of audio tokens into future multimodal language models.